OpenStudy (anonymous):

Powdered aluminum can react with iron (III) oxide in the Thermite reaction to form molten iron and aluminum oxide. Liquid iron is produced because the reaction releases so much energy that the temperature is very high. This liquid can be used to weld railroad rails. What is the balanced molecular equation (bme)? What are the mole ratios? What is the limiting reactant when a mixture of 200.0g of aluminum reacts with 200.0g of iron (III) oxide? What is the mass of liquid iron produced? How many grams of the excess reactant remain after the reaction is complete? In one reaction 135.0g of Fe w

OpenStudy (abmon98):

2Al(s)+Fe2O3(s)--->Fe(l)+Al2O3(s) We begin by determining the number of moles available for each reactant. Number of Moles=Mass(g)/Molar Mass(g/mol) 200/26.98=7.41 moles of Aluminium 200/159.69=1.2524 moles of Iron (III) oxide Use the Mass Method to find the limiting Reagent, we find the masses of iron based on the available mass of each reactant. Look For the Mole Ration 7.41*2 mole of Fe/2 mole of Al2O3*55.845=413.9733 grams of iron 1.2524*2mole of Fe/1 mole of Fe2O3*55.845=139.88grams Clearly Fe2O3 is the limiting Reagent because it produces less aluminium. Mass of Iron Formed is 139,88 gram we Can find the number of moles Aluminium reacted that reacted from the number of moles of Fe2O3 7.41-2.50=4.905 moles of Aluminium remained unreacted, therefore the mass of aluminium unreacted will be 4.905*26.98=132.34 grams I used the equation Number of Moles=Mass(g)/Molar Mass(g/mol)